OpenStudy (anonymous):

What constant do you need to add to each quadratic to complete the square? 1. x^2 - 4x = 0 2. x^2 - 8x = 6 3. 3x^2 + 18x = 24

OpenStudy (amorfide):

the standard formula is \[(x+\frac{ b }{ 2 })^{2} - (\frac{ b }{ 2 })^{2}\] the standard form of a quadratic is \[ax^{2}+bx+c\]

OpenStudy (amorfide):

for x^2 - 4x = 0 b=-4

OpenStudy (anonymous):

2) x^2 -8x -6 +22 =0 x^2-8x+16=0 (x-4)(x-4) =0